OpenStudy (anonymous):

what r the intercepts of 4x^2+16y^2=64

hartnn (hartnn):

@lexigolds9 Hi, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) To get x intercepts put y=0 in 4x^2+16y^2=64 To get y intercepts put x=0 in 4x^2+16y^2=64

OpenStudy (anonymous):

i dont know how to do that. can you tell me the answer? what r the intercepts?

hartnn (hartnn):

i cannot tell you the answers . i will tell you method, u need to get the intercepts on your own, in that way you will learn to get intercepts, ok ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lets hear it.

hartnn (hartnn):

To get x intercepts put y=0 in 4x^2+16y^2=64 what u get ?

OpenStudy (anonymous):

4x^2+16^2+64=0

OpenStudy (anonymous):

??

hartnn (hartnn):

and it would be 4x^2+16^2-64=0 but u did not put y=0 there

OpenStudy (anonymous):

so urs is right and mine is wrong? what next?

hartnn (hartnn):

4x^2+16y^2=64 4x^2 = 64 got this ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so x= ?

OpenStudy (anonymous):

8

hartnn (hartnn):

nopes

OpenStudy (anonymous):

2

hartnn (hartnn):

are these just wild guesses ?

OpenStudy (anonymous):

no i am actually just really stupid at math with variables.

OpenStudy (anonymous):

4 times 2. squared is 64.

OpenStudy (anonymous):

8 times 8.

hartnn (hartnn):

4 x^2 = 64 x^2 = 64/4 x^2 = 16 x= +4 or x=-4 ok?

OpenStudy (anonymous):

ok. it can be -4 or 4?

hartnn (hartnn):

yup, those are x-intercepts. similarly find y-intercepts

OpenStudy (anonymous):

2

OpenStudy (anonymous):

-2

hartnn (hartnn):

that is correct!

OpenStudy (anonymous):

awesome! What next

hartnn (hartnn):

you got both your intercepts

OpenStudy (anonymous):

its asking for intercepts. it says type in an ordered pair. im just confused on which ones. would it be..

OpenStudy (anonymous):

(-4,2)

OpenStudy (anonymous):

it also says there might be more than one ordered pair

hartnn (hartnn):

okk....x-intercepts are of the form (x,0) so it will be (4,0)(-4,0)<----for x-intercepts y-intercepts are of the form (0,y) so it will be (0,2)(0,-2)<----for y-intercepts ok ?

OpenStudy (anonymous):

so there r four different ordered pairs?

hartnn (hartnn):

yup., 2 for x-int. and 2 for y-int.

OpenStudy (anonymous):

thank you!

hartnn (hartnn):

welcome ^_^